OpenStudy (anonymous):

3. Write the equation of the illipse that satisfies the given condition. Sketch each curve. Given problem: Center at (0,0) one vertex at (0,-6),end of a minor axis at (4,0). I need also the sketch please

OpenStudy (anonymous):

ae= 6 b =4 b^2/a^2 = 1-e^2

OpenStudy (anonymous):

find a and b from this and then write the eqn as x^2/a^2 + y^2/b^2 = 1

OpenStudy (anonymous):

so, these are the answer above?

OpenStudy (anonymous):

find a and b and put it in